OpenStudy (anonymous):

How do you solve this question ? Find the derivative y=(lnx)/(x^7)

OpenStudy (anonymous):

The rule is \[\frac{ d }{ dx } \ln|x| = \frac{ 1 }{ x }\]

OpenStudy (anonymous):

what are the next steps?

OpenStudy (anonymous):

Use the Quotient Rule

OpenStudy (anonymous):

I don't think that is the formula I used

OpenStudy (anonymous):

The answer is y=(1-7lnx)/(x^8)

OpenStudy (solomonzelman):

you are not deriving both functions twice in a quotient rule.

OpenStudy (anonymous):

Ohh yes you're right. :$ sorry

OpenStudy (anonymous):

Elaborate please

OpenStudy (solomonzelman):